Specifications for the computer upgrades are 76 minutes and 81 minutes. Estimate the percentage of process output that can be expected to fall within the specifications. (Round your answer to 1 decimal place. Omit the "%" sign in your response.)

SAMPLE
1 2 3 4 5 6
79.2 78.4 79.6 78.9 80.5 79.7
81.9 78.7 79.6 79.8 79.6 80.6
80.0 81.0 78.9 79.7 80.4 79.1
78.4 80.4 80.3 79.4 78.5 80.0
81.0 80.1 80.8 80.6 78.8 81.1


Expected process output            %
Answer #1

Answer: -- According to given data

Lower specification limit = 76

Upper specification limit = 81

So we can see only two values are out of specification limits from 30 observations

Error output = 2/30* 100 = 6.7%

So expected process output = 1 - error output = 100 - 6.7 = 93.3%